# Unified Video Action Model

### v2

@@ Abstract
Figure 1: Unified Video Action Model. (a) UVA features a joint video-action latent representation and decoupled video-action decoding. The joint latent representation effectively models the underlying relationships between video and action sequences, while the decoupled diffusion enables high-speed action inference by bypassing video generation. (b) By leveraging masked training, UVA support flexible input-output combinations for actions and videos. This versatility allows the model to function as a robot policy, video model, forward or inverse dynamics model, or even a combined policy and video planner—all within a unified framework.
@@ I Introduction
We evaluate UVA on seven publicly available benchmarks to assess its diverse capabilities. UVA outperforms or matches state-of-the-art baselines, demonstrating particularly strong performance in multi-task settings. For instance, UVA outperforms the best baseline by 13% in success rate on Libero10 [27] and by 20% on PushT Multitask [9, 14]. The experiments show that UVA can serve as a general-purpose framework for different robotics tasks without compromising performance compared to methods tailored for specific applications. In sum, the unified video action model is:
@@ II Related Work
Video Generation for Policy Learning: Video models aid policy learning by simulating task dynamics and predicting future states. Models like [12, 25] leverage video generation techniques to produce high-quality videos, which are then used for action prediction. The work by [46] leverages video models to generate object flow as an intermediate representation, which captures physical interactions and is used to predict actions for skill transfer across different robotic embodiments and environments. In [21], a video diffusion model is fine-tuned on robotics tasks, with the latent representations from the predicted videos serving as inputs to a policy network for action prediction. Existing approaches that utilize video generation for policy learning often suffer from slow inference speeds, making them impractical for real-world applications. Additionally, these methods often require auxiliary components, such as low-level policies [12] or image-tracking techniques [46], to extract actions from the generated videos. As a result, the final action accuracy suffers from compounded errors in video generation and action prediction.
@@ V-A Simulation Benchmarks
Multi-Task Evaluation: We train one policy for multiple task goals defined by image or text. We introduce a new task, PushT-M, which extends the PushT task to include varying target “T” positions. We evaluate the best-performing checkpoint over 50 rollouts and report its average reward. Libero10 [27] has 10 tasks. We evaluate each task in three different environments with varying random seeds and report the average rewards across all 10 tasks. See Supplementary §X-B for details.
@@ V-B Real-world Benchmarks
  Single-Task $\uparrow$  Multi-Task $\uparrow$  Speed $\downarrow$
 PushT Tool  PushT-M  Libero10
DP-C [9] 0.91 0.95 0.68 0.73 0.50s
DP-T [9] 0.78 0.76 0.63 0.80 0.36s
OpenVLA [22] 0.35 0.18 0.22 0.47 1.52s
UniPi [12] 0.42 0.00 0.19 0.00 24.07s
UVA-action 0.45 0.62 0.46 0.93 0.22s
UVA 0.98 0.88 0.88 0.93 0.23s
@@ V-C Baselines
  • OpenVLA [22] is a state-of-the art vision-language-action built on 7B Llama 2 [38] for multi-task setting. It is trained on a diverse dataset encompassing a wide range of robots, tasks, and environments. We finetune OpenVLA on each task to optimize its performance.
@@ V-D Policy Learning Results
Simulation Multi-Task: Our method is particularly strong in the multi-task setting. Specifically, UVA surpasses the best baseline by 20% on the PushT-M task and by 13% on the Libero10 benchmark. This result demonstrates that UVA model is able to better learn and leverage the general dynamics that are shared across tasks and, therefore, improve overall performance in the multi-task setting.
@@
Our method achieves faster inference speeds by performing diffusion iterations only on the lightweight action head, rather than the entire network as DP-C and DP-T. Thanks to the decoupled design, video generation can be skipped during policy inference, further improving efficiency. For simulation tasks, we use 100 denoise steps for action prediction. For DP-C and DP-T, we follow their original implementations and also perform denoising over 100 steps. With the same number of diffusion steps (Table II), our method achieves faster inference compared to DP-C and DP-T.

### v3

@@ Abstract
Figure 1: Unified Video Action Model. (a) UVA features a joint video-action latent representation and decoupled video-action decoding. The joint latent representation effectively models the underlying relationships between video and action sequences, while the decoupled diffusion enables high-speed action inference by bypassing video generation. (b) By leveraging masked training, UVA supports flexible input-output combinations for actions and videos. This versatility allows the model to function as a robot policy, video model, forward or inverse dynamics model, or even a combined policy and video planner—all within a unified framework.
@@ I Introduction
We evaluate UVA on seven publicly available benchmarks to assess its diverse capabilities. UVA outperforms or matches state-of-the-art approaches, demonstrating particularly strong performance in multi-task settings. For instance, UVA outperforms the best baseline by 20% in success rate on PushT Multitask [10, 16] and by 5% on Libero10 [30]. The experiments show that UVA can serve as a general-purpose framework for different robotics tasks without compromising performance compared to methods tailored for specific applications. In sum, the unified video action model is:
@@ II Related Work
Video Generation for Policy Learning: Video models aid policy learning by simulating task dynamics and predicting future states. Models like [14, 28] leverage video generation techniques to produce high-quality videos, which are then used for action prediction. PAD [19] jointly trains video generation and action prediction; however, it cannot predict future actions independently of future image generation, resulting in slower inference. The work by [49] leverages video models to generate object flow as an intermediate representation, which captures physical interactions and is used to predict actions for skill transfer across different robotic embodiments and environments. In [24], a video diffusion model is fine-tuned on robotics tasks, with the latent representations from the predicted videos serving as inputs to a policy network for action prediction. Existing approaches that utilize video generation for policy learning often suffer from slow inference speeds, making them impractical for real-world applications. Additionally, these methods often require auxiliary components, such as low-level policies [14] or image-tracking techniques [49], to extract actions from the generated videos. As a result, the final action accuracy suffers from compounded errors in video generation and action prediction.
@@ III-B Masked Autoencoder for Observation Prediction
For tasks that involve language instructions, such as Libero10 [30], we incorporate language information using cross-attention in the Transformer. The language input is encoded into a $d$-dimensional token using the CLIP [33] text encoder. To emphasize the language, this token is repeated $M$ times and appended to the $N\times h$ video-action tokens, resulting in a total of $N\times h+M$ tokens. These tokens are then passed through the Transformer to fuse the multimodal information. We take the first $N\times h$ outputs of the Transformer as the joint video-action latent representations, $\{\mathbf{Z}_{t+1},\ldots,\mathbf{Z}_{t+h}\}$.
@@ V UVA as Policy
  Single-Task $\uparrow$  Multi-Task $\uparrow$  Speed $\downarrow$
 PushT Tool  PushT-M  Libero10
DP-C [10] 0.91 0.95 0.68 0.53 0.50s
DP-T [10] 0.78 0.76 0.63 0.58 0.36s
OpenVLA [25] 0.35 0.18 0.22 0.54 1.52s
UniPi [14] 0.42 0.00 0.19 0.00 24.07s
$\pi_{0}$ [2] - - - 0.85 0.09s
$\pi_{0}$-FAST [2] - - - 0.60 0.09s
UVA-action 0.45 0.62 0.46 0.86 0.22s
UVA 0.98 0.88 0.88 0.90 0.23s
@@ V-A Simulation Benchmarks
Multi-Task Evaluation: We train one policy for multiple task goals defined by image or text. We introduce a new task, PushT-M, which extends the PushT task to include varying target “T” positions. We evaluate the best-performing checkpoint over 50 rollouts and report its average reward. Libero10 [30] has 10 tasks. We evaluate each task in 50 different environments with varying random seeds and report the average rewards across all 10 tasks. See Supplementary §X-B for details.
@@ V-C Baselines
  • OpenVLA [25] is a state-of-the art vision-language-action (VLA) built on 7B Llama 2 [41] for multi-task setting. It is trained on a diverse dataset encompassing a wide range of robots, tasks, and environments. We finetune OpenVLA on each task to optimize its performance.
  • $\bm{\pi_{0}}$ [2] is an open-source VLA model designed for general-purpose robot control. It employs a flow matching based architecture to generate continuous action sequences. Trained on a diverse dataset spanning multiple robot embodiments and tasks, $\pi_{0}$ demonstrates strong zero-shot and fine-tuned performance. We use the officially released checkpoints: $\pi_{0}$ and $\pi_{0}$-FAST, both finetuned on Libero tasks.
@@ V-D Policy Learning Results
Simulation Multi-Task: Our method is particularly strong in the multi-task setting. Specifically, UVA surpasses the best baseline by 20% on the PushT-M task and by 5% on the Libero10 benchmark. This result demonstrates that UVA model is able to better learn and leverage the general dynamics that are shared across tasks and, therefore, improve overall performance in the multi-task setting.
Notably, $\pi_{0}$ (3.3B) and $\pi_{0}$-FAST (3.0B) are large models fine-tuned from pretrained checkpoints using the entire Libero dataset, while UVA is smaller (0.5B) and trained only on Libero10. Moreover, $\pi_{0}$ and $\pi_{0}$-FAST use third-person and wrist-view images plus robot proprioception, while UVA relies only on third-person images. Despite its smaller size, limited input, and no external data, UVA outperforms $\pi_{0}$ and $\pi_{0}$-FAST.
@@
Our method achieves faster inference speeds by performing diffusion iterations only on the lightweight action head, rather than the entire network as DP-C and DP-T. Thanks to the decoupled design, video generation can be skipped during policy inference, further improving efficiency. For simulation tasks, we use 100 denoise steps for action prediction. For DP-C and DP-T, we follow their original implementations and also perform denoising over 100 steps. With the same number of diffusion steps (Table II), our method achieves faster inference compared to DP-C and DP-T. Both $\pi_{0}$ and $\pi_{0}$-FAST achieve faster inference speed than other methods.